# Preparation of Papers for IEEE Sponsored Conferences & Symposia*

### v1

@@ II-C Continuous-time Conflict Based Search
CCBS is argued to be a sound and solution-complete MAPFR solver for minimal sum-of-costs [20]. CCBS closely follows CBS [14] by simultaneously building and searching a binary constraint tree (CT).
@@
To explain the CT and how it is searched, we begin by briefly introducing CCBS’s underlying planner Constrained Safe Interval Path Planning (CSIPP) which is covered in more detail in Section V-C2. CSIPP is a variant of SIPP [29] that computes a valid plan for a single agent with minimal duration and that satisfies a given set of constraints $C$. In SIPP, an $A^{*}$ [30] search is performed in the vertex-safe interval state space, with edges in $\mathcal{E}$ providing transitions from one state to the next. In CSIPP, it is instead move actions that provide transitions. Constraints in this context (discussed later in this section) constrain the execution of timed actions. Therefore, CSIPP finds the shortest-duration plan which does not include any timed actions forbidden by constraints in $C$.
@@
In the CT, a node $N$ is associated with a set of constraints $N_{C}$ and a joint plan $N_{\Pi}$. The plans in $N_{\Pi}$ are computed using CSIPP given the constraint set $N_{C}$. Therefore, every plan $\pi_{i}\in N_{\Pi}$ is valid and has minimal duration under the constraints in $N_{C}$. The CCBS search starts at the CT root node $N^{R}$, with constraint set $N^{R}_{C}=\varnothing$. Since $N^{R}_{C}$ is empty, every plan $\pi_{i}\in N^{R}_{\Pi}$ is unconstrained (apart from being valid) and therefore has the shortest possible duration over all valid plans for the corresponding agent $i$. At every iteration of CCBS, a CT node $N$ minimizing $\sigma(N_{\Pi})$ over all unexpanded nodes is selected for expansion. If the joint plan $N_{\Pi}$ is a solution (i.e., conflict-free), then $N_{\Pi}$ is returned as an optimal solution. Otherwise, a branching rule is applied to $N$ to spawn two child nodes. To do so, an arbitrary conflict in $N_{\Pi}$ is selected (say between agents $i$ and $j$) and used to construct a pair of constraints $\langle c_{i},c_{j}\rangle$. These constraints $c_{i}$ and $c_{j}$ respectively constrain the plans of agents $i$ and $j$ to avoid this specific conflict. The child nodes, $N^{i}$ and $N^{j}$, each inherit one of the constraints in addition to all constraints at $N$: $N^{i}_{C}=N_{C}\cup\{c^{i}\}$ and $N^{j}_{C}=N_{C}\cup\{c^{j}\}$. The plans in $N_{\Pi}^{i}$ and $N_{\Pi}^{j}$ are then computed using CSIPP to satisfy $N_{C}^{i}$ and $N_{C}^{j}$, respectively.
@@
We clarify the above by introducing an example in Figure 3, which will later be revisited in more detail. In this example, we consider only two agents $i$ and $j$. At the CT node $N$ (Figure 3(a)) a conflict is detected in the joint plan $N_{\Pi}=\left\{\pi_{i},\pi_{j}\right\}$, illustrated in Figure 3(b). Specifically, $i$ and $j$ collide while respectively executing the timed actions $\langle m^{i}_{2},t^{i}_{2}\rangle$ and $\langle m^{j}_{3},t^{j}_{3}\rangle$. Therefore, $\langle\langle m^{i}_{2},t^{i}_{2}\rangle,\langle m^{j}_{3},t^{j}_{3}\rangle\rangle$ is a conflict. Based on this conflict and a branching rule, a pair of constraints $\langle c^{i},c^{j}\rangle$ is created. We will explore these constraints in more detail later; for now it is sufficient to know that $c^{i}$ constrains when $i$ may execute $m^{i}_{2}$, and $c^{j}$ constrains when $j$ may execute $m^{j}_{3}$, to avoid this specific conflict. Consider the first child node $N^{i}$ (Figure 3(a)) with constraints $N^{i}_{C}=N_{C}\cup\{c^{i}\}$. Using CSIPP, the plans in $N^{i}_{\Pi}$ are computed to satisfy $N^{i}_{C}$. In practice, it is sufficient to only compute $i$’s plan and let all other agents inherit their plans from the parent node, since only $i$ has additional constraints. Figure 3(c) illustrates $N^{i}_{\Pi}$, where agent $i$ now takes a different path through the graph. Similarly, the other child node $N^{j}$ has constraints $N^{j}_{C}=N_{C}\cup\{c^{j}\}$, and a new plan for $j$ is computed using CSIPP. The joint plan $N^{j}_{\Pi}$ is illustrated in Figure 3(d), where $j$ now waits for some time at a vertex before continuing. The resulting joint plans at the child nodes happen to be conflict-free and are therefore both solutions, in this particular example. In general, however, the child nodes’ joint plans may contain conflicts, requiring further CCBS iterations until a solution is found.
(a) The CT node NN and its two children $N^{i}$ and $N^{j}$.
@@
(b) The joint plan $N_{\Pi}=\left\{\pi_{i},\pi_{j}\right\}$ contains a conflict $\langle\langle m^{i}_{2},t^{i}_{2}\rangle,\langle m^{j}_{3},t^{j}_{3}\rangle\rangle$.
(c) The joint plan $N_{\Pi}^{i}$ satisfying $N^{i}_{C}$.
(d) The joint plan $N_{\Pi}^{j}$ satisfying $N^{j}_{C}$.
@@
Figure 3: An example of a branching rule applied to a CT node $N$. Since $N_{\Pi}$ contains a conflict, two child nodes $N^{i}$ and $N^{j}$ are spawned with additional constraints and recomputed plans.
@@
The soundness and solution-completeness proofs of CCBS [20] rely on the CT satisfying the following properties, where $N$ is a parent to $N^{i}$ and $N^{j}$ and $\mathcal{S}(N)$ is the set of all solutions satisfying $N_{C}$:
@@
Property (2) ensures that no solutions are removed when branching a node. At the CT root $N^{R}$, since $N^{R}_{C}=\varnothing$ it holds that all solutions to a MAPFR problem are contained in $\mathcal{S}(N^{R})$. If property (2) holds then $\mathcal{S}(N^{R})=\mathcal{S}(N^{i})\cup\mathcal{S}(N^{j})$ (with $N^{i}$ and $N^{j}$ being children of $N^{R}$) which recursively implies that all solutions are reachable from the CT root. Since the search starts at $N^{R}$, all solutions to a MAPFR problem can be found. Since it is the branching rule that determines $\mathcal{S}(N^{i})$ and $\mathcal{S}(N^{j})$ (by the constraint pair that it constructs), it is the branching rule that determines if property (2) is satisfied.
@@
Property (3) is automatically satisfied by CSIPP. Since plans in $N_{\Pi}$ are computed using CSIPP given constraints $N_{C}$, every plan $\pi_{i}\in N_{\Pi}$ is the shortest-duration valid plan over all valid plans satisfying $N_{C}$. For all nodes $N^{\prime}$ below $N$ we know that $N^{\prime}_{C}\supset N_{C}$ since constraints are only added during a branching, never removed. Adding constraints can never result in shorter-duration plans. Therefore, it holds for all agents $i$ that $\pi_{i,D}\leq\pi_{i,D}^{\prime}$ where $\pi_{i,D}\in N_{\Pi}$ and $\pi_{i,D}^{\prime}\in N_{\Pi}^{\prime}$. Since $\sigma(\Pi)$ is strictly monotonically increasing with the maximum plan duration over all plans $\pi\in\Pi$, it follows that $\sigma(N_{\pi})\leq\sigma(N_{\pi}^{\prime})$, thereby satisfying property (3).
@@
In this article, we focus our attention on the theoretically described branching rule in [20] (the theoretical branching rule) and the publicly available implementation of CCBS [^2] https://github.com/PathPlanning/Continuous-CBS. (the implemented branching rule). Note that details regarding how to implement the theoretical branching rule are provided in [20], however, these details do not match the implemented branching rule and are therefore not considered here.
II-C1 The theoretical branching rule
@@
Given a conflict $\langle\langle a^{i},t^{i}\rangle,\langle a^{j},t^{j}\rangle\rangle\in\pi_{i}\times\pi_{j}$ with $\pi_{i},\pi_{j}\in N_{\Pi}$, the earliest time $t^{i}_{u}>t^{i}$ where agent $i$ executing $\langle a_{i},t^{i}_{u}\rangle$ does not conflict with $j$ executing $\langle a^{j},t^{j}\rangle$ is found. This forms the unsafe interval $t^{i},t^{i}_{u})$. From this, a motion constraint $c_{i}=\langle i,a^{i},[t^{i},t^{i}_{u})\rangle$ is created, specifying that $i$ is forbidden from executing $\langle a^{i},t\rangle$ for any $t\in[t^{i},t^{i}_{u})$. Similarly for $j$, the unsafe interval $t^{j},t^{j}_{u})$ is found, forming the motion constraint $c_{j}=\langle j,a_{j},[t^{j},t^{j}_{u})\rangle$. Therefore the theoretical branching rule uses the constraint pair $\langle\langle i,a_{i},[t^{i},t^{i}_{u})\rangle,\langle j,a_{j},[t^{j},t^{j}_{u})\rangle\rangle$ to branch on $N$. The CT under the theoretical branching rule is shown in [20] to satisfy properties (2) and (3), by which they then conclude that CCBS is sound and solution complete.
@@
Let us now revisit the example in Figure 3, where the conflict $\langle\langle m_{2}^{i},t_{2}^{i}\rangle,\langle m_{3}^{j},t_{3}^{j}\rangle\rangle$ was detected. From this conflict, the motion constraints $c^{i}=\langle i,m_{2}^{i},[t_{2}^{i},t^{i}_{u})\rangle$ and $c^{j}=\langle j,m_{3}^{j},[t_{3}^{j},t^{j}_{u})\rangle$ are formed. At node $N^{i}$ with $N^{i}_{C}=N_{C}\cup\{c^{i}\}$, agent $i$ is no longer permitted to execute $\langle m_{2}^{i},t\rangle$ for $t\in[t_{2}^{i},t^{i}_{u})$. Therefore, an alternative plan is found where $i$ avoids executing $m_{2}^{i}$ entirely. Similarly at node $N^{j}$ where $N^{j}_{C}=N_{C}\cup\{c^{j}\}$, agent $j$ is no longer permitted to execute $\langle m_{3}^{j},t\rangle$ for $t\in[t_{3}^{j},t^{j}_{u})$. Therefore, an alternative plan is found where $j$ waits long enough to eventually execute $m^{j}_{3}$ some time after $t_{u}^{j}$.
II-C2 The implemented branching rule
The implemented branching rule differs from the theoretical implementation by how timed wait actions $\langle w^{j},t^{j}\rangle$ in move-wait conflicts $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ are handled. For such a conflict, the collision interval $I^{c}$ is found. A collision interval is defined as the first contiguous time interval during which $i$ and $j$ are in collision. Then, a vertex constraint $c^{j}=\langle j,v(w^{j}),I^{c}\rangle$ is created, forbidding $j$ from occupying vertex $v(w^{j})$ at any time during $I^{c}$. Besides this, timed move actions in both move-move and move-wait conflicts are handled in the same way as in the theoretical branching rule, resulting in the motion constraint $c^{i}=\langle i,a^{i},[t^{i},t^{i}_{u})\rangle$.

### v2

@@ 2.3 Continuous-time Conflict Based Search
CCBS is argued to be a sound and solution-complete MAPFR solver for minimal sum-of-costs [21]. CCBS closely follows CBS by simultaneously building and searching a binary constraint tree (CT).
@@
To explain the CT and how it is searched, we begin by briefly introducing CCBS’s underlying planner Constrained Safe Interval Path Planning (CSIPP) which is covered in more detail in Section 5.3.2. CSIPP — a variant of Safe Interval Path Planning (SIPP) [30] — computes a minimal-duration valid plan for a single agent that satisfies a given set of constraints $C$. In SIPP, an $A^{*}$ [31] search is performed in the vertex-safe interval state space, with edges in $\mathcal{E}$ providing transitions from one state to the next, and safe intervals defining when an edge may be traversed or a vertex may be occupied. In CSIPP, constraints define the safe intervals and move actions provide transitions between states.
@@
In the CT, a node $N$ is associated with a set of constraints $N_{C}$ and a joint plan $N_{\Pi}$. The plans in $N_{\Pi}$ are computed using CSIPP given the constraint set $N_{C}$, such that every plan $\pi_{i}\in N_{\Pi}$ is valid and has minimal duration under the constraints in $N_{C}$. The CCBS search starts at the CT root node $N^{R}$ with the empty constraint set $N^{R}_{C}=\varnothing$. Since $N^{R}_{C}$ is empty, every plan $\pi_{i}\in N^{R}_{\Pi}$ is unconstrained (apart from being valid) and therefore has the shortest possible duration over all valid plans for agent $i$. At every iteration of CCBS, a node $N$ minimizing $\sigma(N_{\Pi})$ over all unexpanded nodes is selected for expansion. If $N_{\Pi}$ is a solution (i.e., conflict-free), then $N_{\Pi}$ is returned as an optimal solution. Otherwise, a branching rule is applied to $N$ to spawn two child nodes. To do so, an arbitrary conflict in $N_{\Pi}$ is selected (say between agents $i$ and $j$) and used to construct a pair of constraints $\langle c_{i},c_{j}\rangle$. These constraints $c_{i}$ and $c_{j}$ respectively constrain the plans of agents $i$ and $j$ to avoid this specific conflict. The child nodes $N^{i}$ and $N^{j}$ each inherit one of the constraints in addition to all constraints at $N$: $N^{i}_{C}=N_{C}\cup\{c^{i}\}$ and $N^{j}_{C}=N_{C}\cup\{c^{j}\}$. The plans in $N_{\Pi}^{i}$ and $N_{\Pi}^{j}$ are then computed using CSIPP to satisfy $N_{C}^{i}$ and $N_{C}^{j}$, respectively.
@@
We clarify the above by introducing an example in Figure 3, which will later be revisited in more detail. In this example, we consider only two agents $i$ and $j$. At the CT node $N$ (Figure 3(a)) a conflict is detected in the joint plan $N_{\Pi}=\left\{\pi_{i},\pi_{j}\right\}$, illustrated in Figure 3(b). Specifically, $i$ and $j$ collide while respectively executing the timed actions $\langle m^{i}_{2},t^{i}_{2}\rangle$ and $\langle m^{j}_{3},t^{j}_{3}\rangle$. Therefore, $\langle\langle m^{i}_{2},t^{i}_{2}\rangle,\langle m^{j}_{3},t^{j}_{3}\rangle\rangle$ is a conflict. Based on this conflict, a branching rule is used to create a pair of constraints $\langle c^{i},c^{j}\rangle$. We will explore these constraints in more detail later; for now it is sufficient to know that $c^{i}$ constrains when $i$ may execute $m^{i}_{2}$, and $c^{j}$ constrains when $j$ may execute $m^{j}_{3}$, to avoid this specific conflict. Consider the first child node $N^{i}$ (Figure 3(a)) with constraints $N^{i}_{C}=N_{C}\cup\{c^{i}\}$. Using CSIPP, the plans in $N^{i}_{\Pi}$ are computed to satisfy $N^{i}_{C}$. In practice, it is sufficient to only compute $i$’s plan since only the constraints on $i$ have changed, while all other agents can inherit their plans from $N$. Figure 3(c) illustrates $N^{i}_{\Pi}$, where agent $i$ now takes a different path through the graph. Similarly, the other child node $N^{j}$ has constraints $N^{j}_{C}=N_{C}\cup\{c^{j}\}$, and a new plan for $j$ is computed using CSIPP. The joint plan $N^{j}_{\Pi}$ is illustrated in Figure 3(d), where $j$ now waits for some time at a vertex before continuing. In this particular example, the resulting joint plans at the child nodes happen to be conflict-free and are therefore both solutions. In general, however, the child nodes’ joint plans may still contain conflicts, requiring further CCBS iterations until a solution is found.
(a) The CT node $N$ where a conflict is detected in $N_{\Pi}$, and its two children $N^{i}$ and $N^{j}$.
@@
(b) The joint plan $N_{\Pi}=\left\{\pi_{i},\pi_{j}\right\}$ contains a conflict $\langle\langle m^{i}_{2},t^{i}_{2}\rangle,\langle m^{j}_{3},t^{j}_{3}\rangle\rangle$ between agents $i$ and $j$.
(c) The joint plan $N^{i}_{\Pi}$. Additional constraints on agent $i$ are added to $N^{i}_{C}$ to specifically avoid the collision detected at $N$. A new plan $\pi_{i}\in N^{i}_{\Pi}$ satisfying $N^{i}_{C}$ is then computed, where $i$ traverses another path in the graph.
(d) The joint plan $N^{j}_{\Pi}$. Additional constraints on agent $j$ are added to $N^{j}_{C}$ to specifically avoid the collision detected at $N$. A new plan $\pi_{j}\in N^{j}_{\Pi}$ satisfying $N^{j}_{C}$ is then computed, where $j$ waits for some time at a node.
@@
Figure 3: An example of a branching rule applied to a conflict $\langle\langle m^{i}_{2},t^{i}_{2}\rangle,\langle m^{j}_{3},t^{j}_{3}\rangle\rangle$ at CT node $N$. For agent $i$ ($j$), a child node $N^{i}$ ($N^{j}$) is spawned with additional constraints on $i$ ($j$) to avoid this specific conflict. A new plan is then computed which satisfies these constraints.
@@
The soundness and solution completeness proofs of CCBS in [21] rely on the CT satisfying the following properties, where $N$ is a parent to $N^{i}$ and $N^{j}$ and $\mathcal{S}(N)$ is the set of all solutions satisfying $N_{C}$:
@@
Property (2) ensures that no solutions are removed when branching a node. At the CT root $N^{R}$, since $N^{R}_{C}=\varnothing$ it holds that all solutions to a MAPFR problem are contained in $\mathcal{S}(N^{R})$. If property (2) holds then $\mathcal{S}(N^{R})=\mathcal{S}(N^{i})\cup\mathcal{S}(N^{j})$ (with $N^{i}$ and $N^{j}$ being children of $N^{R}$) which recursively implies that all solutions are reachable from the CT root. Since the search starts at $N^{R}$, all solutions to a MAPFR problem are included in the search and therefore can be found. Since it is the branching rule that determines $\mathcal{S}(N^{i})$ and $\mathcal{S}(N^{j})$ (by the constraint pair that it constructs), it is the branching rule that determines if property (2) is satisfied.
@@
Property (3) is automatically satisfied by CSIPP: since plans in $N_{\Pi}$ are computed using CSIPP given constraints $N_{C}$, every plan $\pi_{i}\in N_{\Pi}$ is the shortest-duration valid plan over all valid plans satisfying $N_{C}$. For all nodes $N^{\prime}$ below $N$ we know that $N_{C}\subset N^{\prime}_{C}$ since constraints are only added during a branching, never removed. Adding constraints can never result in shorter-duration plans. Therefore, it holds for all agents $i$ that $\pi_{i,D}\leq\pi_{i,D}^{\prime}$ where $\pi_{i,D}\in N_{\Pi}$ and $\pi_{i,D}^{\prime}\in N_{\Pi}^{\prime}$. Since $\sigma(\Pi)$ is strictly monotonically increasing with the maximum plan duration over all plans $\pi\in\Pi$, it follows that $\sigma(N_{\pi})\leq\sigma(N_{\pi}^{\prime})$, thereby satisfying property (3).
@@
In this article, we focus our attention on two branching rules: the theoretical branching rule (TBR) which is used in the soundness and solution completeness proofs in [21], and the implemented branching rule (IBR) found in the publicly available reference CCBS implementation [^2] https://github.com/PathPlanning/Continuous-CBS..
2.3.1 The Theoretical Branching Rule
@@
The TBR uses the abstract definition of actions, encompassing move and wait actions alike, such that move-move and move-wait conflicts are addressed using the same procedure. Given a conflict $\langle\langle a^{i},t^{i}\rangle,\langle a^{j},t^{j}\rangle\rangle$, the TBR finds the earliest time for each agent to execute its action without colliding with the other agent. That is, the earliest time $t^{i}_{u}>t^{i}$ for agent $i$ to execute $\langle a_{i},t^{i}_{u}\rangle$ without colliding with $j$ executing $\langle a^{j},t^{j}\rangle$ is found. The interval $t^{i},t^{i}_{u})$ is denoted the unsafe interval since $\langle a_{i},t\rangle$ for any $t\in[t^{i},t^{i}_{u})$ certainly conflicts with $\langle a^{j},t^{j}\rangle$. This unsafe interval is used to define the motion constraint $c_{i}=\langle i,a^{i},[t^{i},t^{i}_{u})\rangle$ specifying that $i$ is forbidden from executing $\langle a^{i},t\rangle$ for any $t\in[t^{i},t^{i}_{u})$. Likewise for agent $j$, the unsafe interval $t^{j},t^{j}_{u})$ is found and the motion constraint $c_{j}=\langle j,a_{j},[t^{j},t^{j}_{u})\rangle$ is defined. Letting the conflict be detected at CT node $N$, such that $\langle\langle a^{i},t^{i}\rangle,\langle a^{j},t^{j}\rangle\rangle\in\pi_{i}\times\pi_{j}$ with $\pi_{i},\pi_{j}\in N_{\Pi}$, the TBR uses the constraint pair $\langle c_{i},c_{j}\rangle$ to branch on $N$. The CT using the TBR is shown in [21] to satisfy properties (2) and (3), by which they then conclude that CCBS is sound and solution complete.
@@
Let us now revisit the example in Figure 3 where the conflict $\langle\langle m_{2}^{i},t_{2}^{i}\rangle,\langle m_{3}^{j},t_{3}^{j}\rangle\rangle$ was detected. From this conflict, the motion constraints $c^{i}=\langle i,m_{2}^{i},[t_{2}^{i},t^{i}_{u})\rangle$ and $c^{j}=\langle j,m_{3}^{j},[t_{3}^{j},t^{j}_{u})\rangle$ are formed. At node $N^{i}$ with $N^{i}_{C}=N_{C}\cup\{c^{i}\}$, agent $i$ is no longer permitted to execute $\langle m_{2}^{i},t\rangle$ for $t\in[t_{2}^{i},t^{i}_{u})$. Therefore, an alternative plan is found where $i$ avoids executing $m_{2}^{i}$ entirely. Similarly at node $N^{j}$ where $N^{j}_{C}=N_{C}\cup\{c^{j}\}$, agent $j$ is no longer permitted to execute $\langle m_{3}^{j},t\rangle$ for $t\in[t_{3}^{j},t^{j}_{u})$. Therefore, an alternative plan is found where $j$ waits long enough to eventually execute $m^{j}_{3}$ at $t_{u}^{j}$.
2.3.2 The Implemented Branching Rule
The IBR differs from the TBR by how move-wait conflicts $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ are branched on. To avoid confusion between the IBR and the described implementation in [21], we first describe the latter and then how the IBR differs. With the TBR, the motion constraint $\langle j,w^{j},[t^{j},t^{j}_{u})\rangle$ is created for the move-wait conflict above. Such a motion constraint forbids $j$ from starting the execution of $w^{j}$ at any time in $t^{j},t^{j}_{u})$. That is, $\langle w^{j},t\rangle$ is forbidden for all $t\in[t^{j},t^{j}_{u})$. However, in [21], CSIPP is described to implement such a motion constraint as a vertex constraint $\langle j,v(w^{j}),[t^{j},t^{j}_{u})\rangle$, forbidding $j$ from occupying the vertex $v(w^{j})$ during $t^{j},t^{j}_{u})$. Observe that such a vertex constraint is more constraining than the motion constraint; the motion constraint forbids specifically $w^{j}$ from being executed starting at any time in $t^{j},t^{j}_{u})$, while the vertex constraint forbids every wait action $w$ at the same vertex ($v(w)=v(w^{j})$) from being executed if it overlaps with $t^{j},t^{j}_{u})$, i.e., if $t,t+w_{D}]\cap[t^{j},t^{j}_{u})\neq\varnothing$. Therefore, already here we see a dissonance between the TBR and the described implementation.
Similar to the described implementation, the IBR also implements vertex constraints. However, the IBR uses the intersection interval $\bar{I^{c}}$ instead of the unsafe interval $t^{j},t^{j}_{u})$. Given the conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$, $\bar{I^{c}}$ (formally defined in Definition 3.4) is the time interval during which agent $i$ while executing $\langle m^{i},t^{i}\rangle$ would collide with agent $j$ waiting indefinitely at $v(w^{j})$. Thus, the IBR creates a vertex constraint $\langle j,v(w^{j}),\bar{I^{c}}\rangle$. The intersection interval $\bar{I^{c}}$ is not necessarily the same as the unsafe interval $t^{j},t^{j}_{u})$ since it does not take into account when $j$ begins executing its wait action. For example, if for $\bar{I^{c}}=[\bar{t_{1}^{c}},\bar{t_{2}^{c}})$ we have $\bar{t_{1}^{c}}<t^{j}<\bar{t_{2}^{c}}$, then $j$ can only begin waiting at $v(w^{j})$ from $\bar{t_{2}^{c}}$ else a collision occurs, resulting in the unsafe interval $t^{j},\bar{t_{2}^{c}})\neq\bar{I^{c}}$. Besides this, timed move actions in both move-move and move-wait conflicts are handled in the same way as in the TBR, resulting in the motion constraint $\langle i,a^{i},[t^{i},t^{i}_{u})\rangle$.